I made very rough calibration of the DARM spectra before and after the transition for the second lock on Mar 8.
The cavity pole (expected to be 4.3kHz) was not compensated. Also the servo bump was not compensated.
[Error calibration]
While the DARM/CARM were controlled with ALS, the calibration of them are provided by the ALS phase tracker calibration.
i.e 1 degree = 19.23kHz
This means that the calibration factor is
DARM [deg] * 19.23e3 [Hz/deg] / c [m Hz] * lambda [m] * L_arm [m]
= DARM* 19.23e3/299792458*1064e-9*38.5 = 2.6e-9 *DARM [m]
[Feedback calibration]
Then, the feedback signal was calibrated by the suspension response (f=1Hz, Q=5)
so that the error and feedback signals can match at 100Hz.
This gave me the DC factor of 5e-8.
The spectra at 1109832200 (ALS only, even not on the resonance) and 1109832500 (after DARM/CARM transitions) were taken.
Jenne said that the whitening filters for AS55Q was not on. |